OpenStudy (anonymous):

Can someone please work out the problem 3n-5=7n+11 I don't understand it.

OpenStudy (anonymous):

ok so first combine like terms, you know how to do that correct?

OpenStudy (anonymous):

dear , put n on one sides and numbers on the other side you will be able to solve it iam sure

OpenStudy (anonymous):

4n=-16 .. divide both sides by 4 ... n =-4.

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

if they don't understand the problem and how to do it then just giving them the answer won't help

OpenStudy (anonymous):

you have to carry do you know what you carry over to the other side

OpenStudy (anonymous):

i c .. okay to solve any of this issues . you should do the following .. but the letters at one side and the numbers on the other side with changin the sign of each number or letter you moved .. for example .. iam going to move 7n before the equal it will be -7n .. therfore 3n-7n = -4n .. do you got it ??

OpenStudy (anonymous):

Well, I'm currently failing my algebra honors class. And I took a test on this kind of stuff and got a zero because I missed a whole week of school and my Algebra teacher made me take the test and expects me to make up the whole weeks homework.

OpenStudy (anonymous):

mmm. just kill him :D :D

OpenStudy (anonymous):

i wish...

OpenStudy (anonymous):

did you understand the explanation ??

OpenStudy (anonymous):

kinda. the answers are in the back of the book and it says the answer is -4

OpenStudy (anonymous):

but i dont know how to get that

OpenStudy (anonymous):

yeah sure iam right (h)

OpenStudy (anonymous):

just w8 a secound read carefully the explantion and do what i told u exactly you will get it

OpenStudy (anonymous):

well this is what I did. 3n-5=7n+11 -7n 7n -4n-5=+11

OpenStudy (anonymous):

what do i do now?

OpenStudy (anonymous):

ahhhhhhhh!!!!

OpenStudy (anonymous):

3n-5=7n+11 3n-7n=11+5 right? so then, -4=16 divide both sides by -4 -4 cancels with -4. so u get, 16/-4 =-4 so the answer is -4. Hope I helped :)

OpenStudy (anonymous):

thank youuuuu!

OpenStudy (anonymous):

u're welcome! :D

OpenStudy (anonymous):

sry i got to go yesterday , but i c u got it . snx god :)

OpenStudy (anonymous):

is it with my help or his ??